5. Timeline

Here we document the timeline of changes to SEDM.

Table 1. SEDM Hardware Timeline

Date (UT)

IFU CCD

RC CCD

Description

2021/08/06

05313416

04001312

Adjusted P60 secondary defining points to alleviate vacuum system problem

2021/08/03

05313416

04001312

Repaired Dry Air system online again

2021/07/17

05313416

04001312

P60 secondary issues return

2021/06/23

05313416

04001312

Dry Air system failure

2020/10/29

05313416

04001312

Re-installed repaired Dry Air system and discovered there is no flow control

2020/10/09

05313416

04001312

RC AtoD threshold adjusted so low-background values do not get set to zero

2020/10/06

05313416

04001312

RC serial number fixed, now correct

2020/10/05

05313416

04001312

Adjusted RC alignment and updated RC-IFU offset in pipeline (RC serial still wrong)

2020/10/03

05313416

04001312

Returned to operations with original RC CCD in RC spot, however serial number still shows old IFU CCD number 2803120001 and AtoD threshold too high

2020/10/01

05313416

2803120001

Telescope down for two nights to repair secondary vacuum system and replace RC camera

2020/08/28

05313416

2803120001

Kept original IFU CCD in RC spot, after finding condensation on outside of RC window, disabled dry air system

2020/07/15

05313416

2803120001

Instrument control computer mounted on telescope

2020/07/10

05313416

2803120001

Put original IFU CCD in RC spot due to loose USB port, using spare fibre

2019/09/18

05313416

04001312

P60 secondary axial supports adjusted even closer to retaining ring

2019/09/16

05313416

04001312

P60 primary washed (+12% refl.), RA coupling replaced

2019/09/13

05313416

04001312

P60 secondary axial supports moved to retaining ring to minimize vacuum involvement

2019/09/06

05313416

04001312

P60 secondary re-installed after friction reduction on radial defining points

2019/09/02

05313416

04001312

P60 secondary removed for vacuum system overhaul and testing

2019/08/11

05313416

04001312

S/N measurement star observations taken (focus was good)

2019/06/12

05313416

04001312

First evidence of P60 focus problems

2019/05/20

05313416

04001312

Re-set IFU camera and CCD optics stage: achieved good focus @ 0.4 mm

2019/05/15

05313416

04001312

Laser verification of optical alignment, re-centered IFU CCD

2019/05/14

05313416

04001312

New PI CCD re-installed in IFU position

2019/04/23

2803120001

04001312

Replaced fold mirror with Acton coated fold, moved expander lens, cleaned MLA, Original IFU CCD installed while New PI CCD sent in for board upgrade

2019/04/18

05313416

04001312

Replaced pickoff mirror with Acton coated pickoff, trimmed RC standoffs

2019/02/08

05313416

04001312

Installed instrument with upgraded enclosure, new pickoff/fold mirrors, original optics

UPGRADE #2: Re-did interface plate, enclosure, cable management, new pickoff/fold mirrors, kept original optics.

2019/01/05

05313416

2803120001

Removed instrument for upgrade #2

2018/09/27

05313416

2803120001

Put original IFU CCD in RC spot

2018/08/15

05313416

04001312

S/N measurement star observations taken

2018/08/13

05313416

04001312

Last use of kpy pipeline

2018/07/08

05313416

04001312

Start of pysedm pipeline

2018/06/30

05313416

04001312

Last A/B exposures taken

2018/06/19

05313416

04001312

Installed Jenoptik MLA PN: 107578 520 micron pitch

2018/05/08

05313416

04001312

New PI CCD in IFU spot with corrected SER_NO

2018/01/11

2803120001

04001312

Original RC CCD, New PI CCD sent for repair

2017/09/09

2803120001

08646C7F-

Installed instrument with upgraded enclosure, original optics

UPGRADE #1: Kept mountings and interface plate, re-did enclosure to put electronics outside optics cavity, re-did RC mounting.

2017/02/28

2803120001

08646C7F-

Removed instrument for upgrade #1

2016/07/07

2803120001

08646C7F-

New PI CCD in RC spot with incorrect SER_NO (should be 05313416)

2016/06/02

2803120001

04001312

Original RC CCD (now repaired)

2016/04/25

2803120001

08098513

Original IFU CCD (now repaired)

2016/03/26

none

08098513

Moved loaner IFU CCD to RC

2016/02/29

08098513

04001312

Loaner CCD from PI for IFU

2015/11/15

2803120001

04001312

Earliest data set in /scr2/sedm/raw on pharos

2014/01/31

2803120001

04001312

Earliest definitive appearance of ‘stripes’ in original IFU CCD data

2013/03/12

2803120001

04001312

Earliest data set in /scr2/npk/raw on pharos

Table 1 shows the dates when major changes to the instrument hardware were made. This table does not capture adjustments of optics or minor hardware changes. Column 1 is the date the change was made, column 2 is the IFU CCD serial number, column 3 is the RC CCD serial number, and column 4 describes the hardware change.

Last updated on 11 August 2021